OpenStudy (anonymous):

I need some help with this: to be calculated as a system: x+y=4 xy=3

OpenStudy (anonymous):

it need to be calculated as R x R

mathslover (mathslover):

x + y = 4 -------------------- (1) xy = 3 -------------------- (2) These two equations are given to you.

mathslover (mathslover):

square the equation (1) both sides.

mathslover (mathslover):

What do you get?

mathslover (mathslover):

@paperboy please follow the above instruction and let me know what you got...

hero (hero):

What about using substitution?

OpenStudy (anonymous):

@mathslover my main language is not romanian so corect me if I'm wrong x^2 + y^2 = 16

hero (hero):

y = 4 - x x(4 - x) = 3

mathslover (mathslover):

@paperboy - no , (x+y)^2 = x^2 + y^2 + 2xy and yes @Hero 's method may be a lot easier.

OpenStudy (wolf1728):

How about solving for 'y' in the equation xy=3 ? y = 3/x then substituting this into the other equation? x + y =4 x + 3/x = 4 Multiplying both sides by 'x' x² +3 = 4x x² -4x +3 =0 You should be able to finish this calculation.

mathslover (mathslover):

That may work too @Hero But the method I am using here is the one I love the most! :) It will go like this : x + y = 4 Squaring both sides : (x+y)^2 = 16 x^2 + y^2 + 2xy = 16 x^2 + y^2 + 2(3) = 16 [as xy = 3] x^2 + y^2 = 10 (x-y)^2 = x^2 + y^2 - 2xy = 10 - 2(3) = 4 therefore , (x-y)^2 = 4 or x-y = 2 thus, x + y = 4 and x - y = 2 Thus, 2x = 6 , x = 3 and y = 1

hero (hero):

You only do that if you want to intentionally make the steps longer. There's no need to square both sides.

hartnn (hartnn):

i would have let mathslover finish explaining his method, then jump in and say, you can do it this way too.... interrupting in between will just lead to confusion.....just saying

hero (hero):

@hartnn, no one interrupted anyone

mathslover (mathslover):

Well, hero, I squared both sides to get the value of x^2 + y^2 which i can put in (x-y)^2 expression and solve for x-y. Let us stop our discussion and concentrate on teaching @paperboy in proper way. Hero, your method is a bit easier, so you please teach him your method.

hero (hero):

No, you can continue with your method. I'm just going to leave now since I'm interrupting. Thanks @hartnn

mathslover (mathslover):

As you wish, @paperboy : First of all, (x+y)^2 = x^2 + y^2 + 2xy not x^2 + y^2 So, you get now : (x^2 + y^2 + 2xy) = 16 right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

so now, can you substitute the value of xy there in the equation : x^2 + y^2 + 2xy = 16

OpenStudy (anonymous):

but i need to solve it as R x R ( R*R) and i think that @wolf1728 's method was what the exercise asked for

mathslover (mathslover):

Oh, didn't notice that. Sorry, then use @Hero 's method... or @wolf1728 's method.

mathslover (mathslover):

Hero may like to teach that method now!

hero (hero):

No, I just came to say good job on your method @mathslover. It's a good alternative method.

mathslover (mathslover):

Thanks Hero. Paper boy, you are given with xy = 3 and since x + y = 4 or y = 4 - x therefore : xy = 3 becomes : x(4-x)= 3

mathslover (mathslover):

can u solve it now?

OpenStudy (anonymous):

I think I can, thanks

mathslover (mathslover):

Great, good luck! Also, As you are new to this site, \(\Huge\tt \color{#CC0000}W\color{#CD0B00}e\color{#CF1700}l \color{#D12200}c\color{#D32E00}o\color{#D53900}m\color{#D74500}e~\color{#DB5C00}t\color{#DD6800}o~\color{#E17F00}O\color{#E38B00}p\color{#E59600}e\color{#E7A200}n\color{#E9AD00}S\color{#EBB900}t\color{#EDC500}u\color{#EFD000}d\color{#F1DC00}y\color{#F3E700}!\color{#F5F300}!\) Good Luck for your learning on this site. Do let me know if you get any problem while working on this site.

OpenStudy (anonymous):

@mathslover can you help solve this as you last said? It seems like I'm not doing it right "you are given with xy = 3 and since x + y = 4 or y = 4 - x therefore : xy = 3 becomes : x(4-x)= 3"

mathslover (mathslover):

Great to see that you tried it yourself first. Nice spirit. See, x(4-x)=3 4x - x^2 = 3 right?

OpenStudy (anonymous):

yes, now what i do next?

mathslover (mathslover):

I can write it as : x^2 + 3 - 4x = 0

mathslover (mathslover):

or x^2 - 4x + 3 = 0 or x^2 - 3x - x + 3 = 0 x(x-3) - (x-3) = 0 (x-1)(x-3) = 0 so, x = 1 or x = 3

mathslover (mathslover):

getting it?

OpenStudy (anonymous):

yes, thanks a lot